OpenStudy (anonymous):

whats the volume of the solid obtained when rotating the region under the graph of the function f(x)= x^(13/12)?? on the interval [1,3]??? need help

OpenStudy (anonymous):

vol = INT pi f(x)^2 dx between x = 1 and 3 = INT pi x^169/144 dx = pi [x^313/144 * 144/313] btw 1 and 3 = = pi [3^313/144 * 144/313 -1^313/144 * 144/313 ]

OpenStudy (anonymous):

OK - can u follow it ?? its an awkward one as u have to square x^(13/12)

OpenStudy (anonymous):

ya see i already had all that but its telling me i got the wrong answer

OpenStudy (anonymous):

the vol of revolution = INT pi f(x)^2 dx

OpenStudy (anonymous):

did u make a mistake in the calculation perhaps what was the value the book gave?

OpenStudy (anonymous):

(24.573pi/13

OpenStudy (anonymous):

ok i'll work out what i've got

OpenStudy (anonymous):

ya thats not even the right answer either

OpenStudy (anonymous):

well i'm pretty sure the integration is correct but i've got to go - sorry creeg - hope u get good answer - i'll check back in an hour to see if u still want help

OpenStudy (anonymous):

ill be here

OpenStudy (anonymous):

ok - i'm back and i can see i messed up on squaring x^ 13/12 its not x^169/144 you just multiply the power by 2 so x ^(13/12) squared = x ^ (26/12) = x ^ (13/6) so the working is:- INT pi x^13/6 dx = pi [x^19/6 * 6/19] btw 1 and 3 = = pi [3^19/6 * 6/19 -1^19/6 * 6/19] using my calculator I get =9.924 pi